OpenStudy (anonymous):

What is the exact value of 7 times q to the fourth power all over the square root of q to the fifteenth power.? Simplify if possible. 7 the square root of q all over q to the fourth power. 7 the square root of q all over q cubed. 7 the square root of q all over q to the sixth power. 7 the square root of q all over q to the fifth power.

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

so you have 7q^4 ------ \(\sqrt{q^15}\)

OpenStudy (anonymous):

YA

OpenStudy (ivettef365):

you need to multiply by \(\sqrt{q^15}\) so you have 7q^4 \(\sqrt{q^15}\) ----- . --------- \(\sqrt{q^15}\) \(\sqrt{q^15}\)

OpenStudy (anonymous):

OK

OpenStudy (ivettef365):

so you have 7q^4\(\sqrt{q15}\) ---------------- q^15 when you multiply a square root by another square root you can take out the number

OpenStudy (ivettef365):

now it could be solved as follows 7q^4 . q^7 \(\sqrt{q}\) \(\sqrt{q15}\) = \(\sqrt{q7.q7.q}\), so take out q^7 --------------- q^15

OpenStudy (anonymous):

OK

OpenStudy (ivettef365):

that is equal to 7q^11\(\sqrt{q}\) -------------- q^15

OpenStudy (ivettef365):

now you can simplify the q^11 with the q^15 and you have left 7\(\sqrt{q}\) ------ q^4

OpenStudy (anonymous):

THANK YOU What is the product of 6the square root of the quantity 12 times k. • 4kthe square root of 3.? Simplify if possible. 72kthe square root of k. 72kthe square root of the quantity 2 times k. 144kthe square root of k. 144kthe square root of the quantity 2 times k.

OpenStudy (anonymous):

CAN YOU STILL HELP ME